In order to buy a house you want to accumulate a down payment of $15,000. You can do that by putting a certain sum of money every month in a savings account for the next four years. The account credits interest at the rate of six percent. What is your required monthly deposit?

Relevant Periods=

Relevant Rate=

Deposit Amount=
Answer #1

Amount to be accumulated = $15,000

Interest rate = 6%

The deposits has to be made monthly. So, interest rate has to be divided by 12.

Relevant rate = 6/12 = 0.5%

Time period = 4 years

The deposit has to be made monthly. So, time period has to be multiplied by 12.

Relevant period = 4 * 12 = 48

Calculate the monthly deposit -

Monthly deposit = Amount to be accumulated (A/P, i, n)

Monthly deposit = $15,000 (A/P, 0.5%, 48)

Monthly deposit = $15,000 * 0.02349 = $352.35

Thus,

The required monthly deposit is $352.35